Select the CORRECT statements.

Select the CORRECT statements.


Given the variance of 5.78, the standard deviation is 7.73.


The coefficient of variation is equal to 150% when the mean is 8 and the standard deviation is 12.


Prices are the example of the nominal scale of measurement.


Given a normal (bell-shaped) distribution with the mean of 12 and the standard deviation of 4, approximately 2.5% of observations will be above 20.


It is usually good idea to have more than 35 classes in frequency table (or histogram).
Answer #1

Ans:
1. Correct
Coefficient of variation = S.D/mean = 12/8 = 1.5 or 150%


2. Wrong
Generally making of 6 to 15 number of classes is a good idea for a frequency table


3. Correct
Mean = 12
S.D = 4
X = 20
z = (X-Mean)/S.D = (20 - 12)/4 = 2
P(X>20) = 2.5% (From standard normal table)


4. Wrong
Prices are the ratio scale of data


5. Wrong
Variance = (S.D)^2
But here the above condition is not satisfied


answered by: ANURANJAN SARSAM